OpenStudy (anonymous):

simplify this radical expression sqrt 121y^10

OpenStudy (anonymous):

\[\sqrt{121*y^{10}}=\sqrt{11^2*(y^{5})^2}=11\left| y^5 \right|\]

OpenStudy (anonymous):

11y^5, 11*11 = 121 y^10 = (y^2)^5 so the square root of all of that is 11y^5.

OpenStudy (anonymous):

thanks. ;]